OpenStudy (anonymous):

i know the answer to this problem but need someone to explain it to me 3[2x-(x+7)]=5(x-3) solve X.

OpenStudy (radar):

I think this problem is to determine if you know the order of operations. step 1. Remove the inner parenthesis. 3[2x-x-7]=5x-15 remove brackets 6x-3x-21=5x-15 combining similar terms 3x-21=5x-15 adding 21 and subtracting 5x from both sides -2x=6 dividing both sides by =2 x=-3 Is this what you got?

OpenStudy (anonymous):

yes thats the answer on the key! thank u so much!

OpenStudy (radar):

*-2 instead of =2 I am correcting the last line

OpenStudy (radar):

it should read: -2x=6 dividing both sides by -2

OpenStudy (radar):

Good luck with your studies.

OpenStudy (anonymous):

thanks i'll need it TEST TOMORROW!

OpenStudy (radar):

You're getting ready, just relax and you will do O.K. tomorrow.